MHT CET 2024 Mathematics question
The bulbs are drawn successively with replacement from a lot containing 10% defective bulbs then the probability that there is at least one defective bulb is
Answer: (C)
Solution
Each draw (with replacement) is an independent Bernoulli trial with and . Taking draws as indicated by the options, $P(at least one defective)=1-P(\text{none
See the full step-by-step solution
Create a free account to read the complete working, and practise questions like this in a timed test.
Create a free account